Find p(O), p(l ) and p(2) for each of the following
polynomials:
(J) p(y) = y2 -'+ 1 (i) p(t) = 2 + t + 2tf - t3
(iii) p(x) • (iv) p(x) = (x.- l)(x + 1).

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
i) p(O) = 0 - 0 + 1 = 1;p(l) = 1- 1+ 1= l; p(2) = 4 - 2 + 1= 3. (u) p(O) = 2 + 0 + 0 - 0 = 2; p(l) = 2 + 1+ 2 - 1= 4; p(2) - 2 + 2 + 8 - 8 • 4. (iii) p(O) = O; p(l) = 1; p(2) = 8.